# FinTagging: An LLM-ready Benchmark for Extracting and Structuring Financial Information

### v1

@@ Appendix A Limitations
In this work, we propose FinTagging and conduct a comprehensive analysis of LLM performance on the task. However, several limitations remain. First, we have not yet evaluated a broader range of models, including recent releases such as GPT-4.1, LLaMA-4, and Qwen-3. We are currently conducting these experiments and will report the results in future work. Second, although we provide detailed empirical results, we did not perform statistical significance testing, as evaluating the complete benchmark, especially for models like GPT and DeepSeek, requires substantial time. This analysis is also in progress and will be included in future updates. Third, we collected only a small dataset for fine-tuning PLMs and did not construct task-specific training data for LLMs. Future work may explore building LLM-based financial tagging agents through targeted fine-tuning to further enhance performance.
Appendix B Literature Review
@@
B.1 XBRL Tagging benchmark
@@
FiNER systematically benchmarked several neural architectures on the finer-139 dataset to address numeric-heavy XBRL tagging [14]. The initial experiments showed that standard BERT underperforms due to subword fragmentation, then the authors introduced pseudo-token strategies replacing numerals with [NUM] or [SHAPE] tokens to stabilize label assignment across fragmented numeric spans. These strategies, combined with domain-specific pretraining on SEC-BERT, significantly improved tagging performance, reaching 82.1 micro-F1 without the need for computationally expensive CRF layers. Their experiments demonstrated that subword-aware models with numeric-aware pseudo-tokens outperform both word-level BiLSTMs and vanilla BERT, particularly in numeric-heavy contexts, and avoid nonsensical label sequences. FNXL extended this benchmarking paradigm to a much larger label space of 2,794 US-GAAP tags, reframing the task as an extreme classification problem [23]. They compared the FiNER sequence labeling approach with a two-step pipeline that first identifies numeric spans and then assigns labels using AttentionXML. While FiNER achieved stronger micro-F1 (75.84), reflecting better performance on frequent tags, AttentionXML outperformed FiNER in macro-F1 (47.54), highlighting its strength in predicting infrequent, tail-end labels. FNXL further evaluated both models under a Hits@k setting, confirming that label recommendations from the AttentionXML pipeline could substantially reduce manual effort and maintain high inter-annotator agreement. Together, these benchmarks reveal the need for context-aware reasoning and label-ranking mechanisms in realistic XBRL tagging scenarios.
B.2 XBRL Tagging Methods
@@
The previous studies also explored the approaches address scalability, semantic ambiguity, and reasoning gaps in XBRL tagging to improve the performance. Saini et al. [21] proposed GalaXC is a graph-based extreme classification framework that jointly learns over document-label graphs with per-label attention across multi-hop neighborhoods. By integrating label metadata and transitive label correlations, GalaXC outperformed leading deep classifiers by up to 18% in micro-F1 on standard benchmarks and achieved 25% gains in warm-start scenarios where partial labels are available. Moreover, Wang et al. [28] addressed the practical challenge of custom tag standardization through a semantic similarity pipeline that leverages TF-IDF, Word2Vec, and FinBERT embeddings. Although unsupervised, the method was tested across nearly 200,000 custom tags from SEC filings between 2009 and 2022, and showed strong alignment performance, with vector-based mappings identifying viable standard tag candidates for a substantial proportion of non-compliant elements—offering a low-cost, interpretable solution for downstream financial analysis. Shifting focus from classification to comprehension, XBRL-Agent evaluated the capabilities of large language models to reason over full XBRL reports [10]. The authors introduced two task types—domain taxonomy understanding and numeric reasoning and found that base LLMs often hallucinated or misinterpreted financial content. To overcome these issues, XBRL-Agent incorporated retrieval-augmented generation (RAG) and symbolic calculators within an LLM-agent framework. The enhanced system achieved a 17% accuracy gain on domain query tasks and a 42% boost on numeric reasoning queries compared to base LLMs, validating the utility of modular tool augmentation. These improvements enabled reliable multi-step reasoning over complex disclosures such as debt instruments and derivative gains, which are difficult to capture using span-level classifiers. Collectively, these works broaden the methodological landscape of XBRL tagging from graph-based label propagation and embedding-based normalization to LLM-driven report analysis and point to a hybrid future where structured priors and reasoning tools jointly support accurate, scalable financial information extraction.
B.3 Financial Evaluation Benchmarks
@@
In parallel to XBRL-specific advances, the financial NLP community has developed comprehensive benchmarks to assess broader capabilities in information extraction, numerical reasoning, and document understanding. FiNER-ORD [22] introduced a high-quality, domain-specific NER dataset annotated over financial news, emphasizing general entity types like persons, organizations, and locations. While not numerically focused like FiNER-139, it highlights the lexical diversity of financial discourse and establishes a strong baseline for testing pretrained and zero-shot LLMs in real-world financial NER scenarios. FinQA [4] pushed toward explainable QA by pairing expert-written questions with annotated multi-step reasoning programs derived from earnings reports. ConvFinQA [5] extended this challenge to conversational contexts, simulating real-world question flows over sequential financial queries. TAT-QA [35] focused on hybrid tabular-text reasoning and required models to align cell values and document narratives, often involving aggregation, comparison, and unit-scale interpretation. Pixiu [33] introduced a broader evaluation framework by releasing FinMA, a financial LLM instruction-tuned across five tasks, and assessing it on a new benchmark covering sentiment classification, QA, summarization, NER, and stock prediction. BizBench [12] framed financial QA as program synthesis over realistic, multi-modal contexts, integrating reasoning, code generation, and domain knowledge into a single evaluation pyramid. While these benchmarks highlight the growing ability of models to integrate structured and unstructured financial data, they overlook taxonomy-driven fact alignment and do not support the structured output formats required for XBRL tagging.
Appendix C The Statistics of the Company
Table 9 lists the 30 publicly traded companies whose 2024 annual 10-K filings were used in our study. For each company, we report its Central Index Key (CIK), legal name, stock ticker symbol, and the fiscal year-end filing date. These reports were retrieved from the SEC EDGAR system and conform to the XBRL standard. The selected filings span a diverse set of industries and ensure a representative sample for analyzing structured and unstructured financial disclosures. This company list complements the dataset statistics summarized in Table 2.
Table 9: Company filing summary (2024) for benchmark data
CIK Company Name Ticker Filing Date
0001163739 NABORS INDUSTRIES LTD NBR 2024-12-31
0001418819 Iridium Communications Inc. IRDM 2024-12-31
0001993004 NorthWestern Energy Group, Inc. NWE 2024-12-31
0000024741 CORNING INC /NY GLW 2024-12-31
0000079282 BROWN & BROWN, INC BRO 2024-12-31
0001364479 HERC HOLDINGS INC HRI 2024-12-31
0000936340 DTE ENERGY CO DTM 2024-12-31
0000085535 ROYAL GOLD INC RGLD 2024-12-31
0000086312 TRAVELERS COMPANIES, INC. TRV 2024-12-31
0001968915 PHINIA INC. PHIN 2024-12-31
0000093751 STATE STREET CORP STT 2024-12-31
0000106640 WHIRLPOOL CORP WHR 2024-12-31
0000888491 OMEGA HEALTHCARE INVESTORS INC OHI 2024-12-31
0001616862 Axalta Coating Systems Ltd. AXTA 2024-12-31
0000922224 PPL Corp PPL 2024-12-31
0000066756 ALLETE INC ALE 2024-12-31
0001411207 Allison Transmission Holdings Inc ALSN 2024-12-31
0001478242 IQVIA HOLDINGS INC. IQV 2024-12-31
0001544400 NFiniTi inc. NFTN 2024-12-31
0000851205 COGNEX CORP CGNX 2024-12-31
0001318220 Waste Connections, Inc. WCN 2024-12-31
0000048898 HUBBELL INC HUBB 2024-12-31
0000937098 TRINET GROUP, INC. TNET 2024-12-31
0000910108 LXP Industrial Trust LXP 2024-12-31
0001932393 GE HealthCare Technologies Inc. GEHC 2024-12-31
0001770787 10x Genomics, Inc. TXG 2024-12-31
0001026214 FEDERAL HOME LOAN MORTGAGE CORP FMCC 2024-12-31
0000004904 AMERICAN ELECTRIC POWER CO INC AEP 2024-12-31
0001637459 Kraft Heinz Co KHC 2024-12-31
0001713445 Reddit, Inc. RDDT 2024-12-31
Appendix D Evaluation Metrics
@@
To provide a fair evaluation of overall benchmark performance, we adopt a set of metrics, focusing primarily on macro-level and micro-level evaluation strategies inspired by the previous work [23]. Macro-level evaluation computes precision, recall, and F1 scores independently for each BIO-concept label derived from the US-GAAP taxonomy, and then averages them without weighting. This ensures that each concept, including rare or infrequent ones, contributes equally to the final score—making it especially suitable for domains with skewed label distributions. In contrast, micro-level evaluation aggregates token-level true positives, false positives, and false negatives across all labels before computing precision, recall, and F1. This approach emphasizes the model’s overall tagging accuracy by treating every token equally, and thus better reflects performance on frequent concepts. Together, these two metrics provide a balanced view of both per-concept performance and overall tagging quality.
[…]
$\text{Precision}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{P}|}$ (5)
$\text{Recall}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{G}|}$ (6)
[…]
@@
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\hat{c}_{i},c_{i}^{*})$ (8)
where $\delta(\hat{c}_{i},c_{i}^{*})=1$ if $\hat{c}_{i}=c_{i}^{*}$, and 00 otherwise.
Appendix E Evaluation Models Details

### v2

@@ Appendix A Limitations
In this work, we propose 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and conduct a comprehensive analysis of LLM performance on the task. However, several limitations remain. First, we have not yet evaluated the latest models, such as GPT-5, gpt-oss, and Qwen3-235B. We are currently conducting these experiments and will report the results in future work. Second, we collected only a small dataset for fine-tuning PLMs and did not construct task-specific training data for LLMs. Future work may explore building LLM-based financial tagging agents through targeted fine-tuning to further enhance performance.
Appendix B Significance analysis
To further examine the comparative performance across models, we conducted a pairwise significance analysis, with the results summarized in Figure 4. The upper-triangular matrix in SubFigure 4(a) presents the pairwise significance outcomes, where a value of 1 indicates that the row model’s performance is significantly different from the column model. In contrast, 0 indicates that the difference is not statistically significant. The diagonal elements are set to zero, since self-comparisons are not meaningful.
(a) Model-vs-Model significance matrix.
(b) Per-model count of significant pairwise differences.
Figure 4. Visualization of significance analysis across models. (a) shows the pairwise-bootstrap significance matrix, while (b) summarizes how many significant differences each model has relative to others.
As shown in SubFigure 4(b), the number of significant pairwise differences varies across models, reflecting heterogeneous performance behaviors on the evaluated benchmark. The models DeepSeek-R1-Distill-Qwen-32B and Qwen3-32B exhibit the largest numbers of significant differences, followed by GPT-4o, Llama-4, and Qwen3-14B, suggesting that their performance diverges more noticeably from other models. These results may indicate that models within the Qwen family display stronger variability across comparisons, potentially due to differences in training objectives or architectural configurations. In contrast, Deepseek-V3, Llama-3.3-70B-Instruct, gemma-2-27b-it, and Fino1-8B demonstrate relatively fewer significant differences, implying that their performance remains more stable and comparable to other high-performing models. Overall, these findings suggest that larger or instruction-tuned models tend to produce more consistent outcomes, while models with smaller sizes or domain-oriented tuning show greater variation and clearer statistical separations under significance testing.
Overall, this analysis highlights not only which models outperform or underperform others but also which models behave similarly when subjected to statistical evaluation. The pairwise significance matrix, therefore, provides an interpretable view of model robustness and performance distinctiveness beyond average accuracy metrics.
Appendix C Literature Review
@@
C.1. XBRL Tagging benchmark
@@
FiNER systematically benchmarked several neural architectures on the finer-139 dataset to address numeric-heavy XBRL tagging (Loukas et al., 2022). The initial experiments showed that standard BERT underperforms due to subword fragmentation, then the authors introduced pseudo-token strategies replacing numerals with [NUM] or [SHAPE] tokens to stabilize label assignment across fragmented numeric spans. These strategies, combined with domain-specific pretraining on SEC-BERT, significantly improved tagging performance, reaching 82.1 micro-F1 without the need for computationally expensive CRF layers. Their experiments demonstrated that subword-aware models with numeric-aware pseudo-tokens outperform both word-level BiLSTMs and vanilla BERT, particularly in numeric-heavy contexts, and avoid nonsensical label sequences. FNXL extended this benchmarking paradigm to a much larger label space of 2,794 US-GAAP tags, reframing the task as an extreme classification problem (Sharma et al., 2023). They compared the FiNER sequence labeling approach with a two-step pipeline that first identifies numeric spans and then assigns labels using AttentionXML. While FiNER achieved stronger micro-F1 (75.84), reflecting better performance on frequent tags, AttentionXML outperformed FiNER in macro-F1 (47.54), highlighting its strength in predicting infrequent, tail-end labels. FNXL further evaluated both models under a Hits@k setting, confirming that label recommendations from the AttentionXML pipeline could substantially reduce manual effort and maintain high inter-annotator agreement. Together, these benchmarks reveal the need for context-aware reasoning and label-ranking mechanisms in realistic XBRL tagging scenarios.
C.2. XBRL Tagging Methods
@@
The previous studies also explored the approaches address scalability, semantic ambiguity, and reasoning gaps in XBRL tagging to improve the performance. Saini et al. (Saini et al., 2021) proposed GalaXC is a graph-based extreme classification framework that jointly learns over document-label graphs with per-label attention across multi-hop neighborhoods. By integrating label metadata and transitive label correlations, GalaXC outperformed leading deep classifiers by up to 18% in micro-F1 on standard benchmarks and achieved 25% gains in warm-start scenarios where partial labels are available. Moreover, Wang et al. (Wang, 2023) addressed the practical challenge of custom tag standardization through a semantic similarity pipeline that leverages TF-IDF, Word2Vec, and FinBERT embeddings. Although unsupervised, the method was tested across nearly 200,000 custom tags from SEC filings between 2009 and 2022, and showed strong alignment performance, with vector-based mappings identifying viable standard tag candidates for a substantial proportion of non-compliant elements—offering a low-cost, interpretable solution for downstream financial analysis. Shifting focus from classification to comprehension, XBRL-Agent evaluated the capabilities of large language models to reason over full XBRL reports (Han et al., 2024). The authors introduced two task types—domain taxonomy understanding and numeric reasoning and found that base LLMs often hallucinated or misinterpreted financial content. To overcome these issues, XBRL-Agent incorporated retrieval-augmented generation (RAG) and symbolic calculators within an LLM-agent framework. The enhanced system achieved a 17% accuracy gain on domain query tasks and a 42% boost on numeric reasoning queries compared to base LLMs, validating the utility of modular tool augmentation. These improvements enabled reliable multi-step reasoning over complex disclosures such as debt instruments and derivative gains, which are difficult to capture using span-level classifiers. Collectively, these works broaden the methodological landscape of XBRL tagging from graph-based label propagation and embedding-based normalization to LLM-driven report analysis and point to a hybrid future where structured priors and reasoning tools jointly support accurate, scalable financial information extraction.
C.3. Financial Evaluation Benchmarks
@@
In parallel to XBRL-specific advances, the financial NLP community has developed comprehensive benchmarks to assess broader capabilities in information extraction, numerical reasoning, and document understanding. FiNER-ORD (Shah et al., 2023) introduced a high-quality, domain-specific NER dataset annotated over financial news, emphasizing general entity types like persons, organizations, and locations. While not numerically focused like FiNER-139, it highlights the lexical diversity of financial discourse and establishes a strong baseline for testing pretrained and zero-shot LLMs in real-world financial NER scenarios. FinQA (Chen et al., 2021) pushed toward explainable QA by pairing expert-written questions with annotated multi-step reasoning programs derived from earnings reports. ConvFinQA (Chen et al., 2022) extended this challenge to conversational contexts, simulating real-world question flows over sequential financial queries. TAT-QA (Zhu et al., 2021) focused on hybrid tabular-text reasoning and required models to align cell values and document narratives, often involving aggregation, comparison, and unit-scale interpretation. Pixiu (Xie et al., 2023) introduced a broader evaluation framework by releasing FinMA, a financial LLM instruction-tuned across five tasks, and assessing it on a new benchmark covering sentiment classification, QA, summarization, NER, and stock prediction. BizBench (Koncel-Kedziorski et al., 2023) framed financial QA as program synthesis over realistic, multi-modal contexts, integrating reasoning, code generation, and domain knowledge into a single evaluation pyramid. While these benchmarks highlight the growing ability of models to integrate structured and unstructured financial data, they overlook taxonomy-driven fact alignment and do not support the structured output formats required for XBRL tagging.
Appendix D Data and Ticker Information
D.1. Overall Data Annotation
Different from the filtered dataset described in Section 3.3, we construct an overall annotation dataset based directly on the unfiltered raw data. As summarized in Table 11, the dataset consists of both sentence-level and table-level instances drawn from the narrative and tabular content of financial reports. In total, we annotated 15,986 sentence instances and 12,801 table instances, including 7,768 positive and 8,218 negative sentences, as well as 8,709 positive and 4,092 negative tables. On average, textual sentences contain 80.91 tokens, while table sequences are considerably longer, averaging 1,212.42 tokens. Across all data, we identified 261,457 numerical entities linked to 3,953 unique US-GAAP concepts.
Table 11. Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
Item Sentence Table
Positive instances 7,768 8,709
Negative instances 8,218 4,092
Avg. Tokens/S $80.91\pm 63.62$ $1212.42\pm 1421.76$
Avg. Entities/S $1.24\pm 1.82$ $18.87\pm 37.16$
Avg. Concepts/S $1.24\pm 1.82$ $18.87\pm 37.16$
Total Entities 261,457
Unique Concepts 3,953
D.2. The Statistics of the Tickers
Figure 5, Figure 6, and Figure 7 illustrate the distribution of 142 tickers across industry sectors, market capitalization categories, and geographic regions, highlighting the diversity of our collected financial reports. Considering that practical XBRL tagging practices often vary by industry, company size, and legal jurisdiction, we curated a diverse set of companies covering all 11 major industry sectors. The sample maintains a balanced distribution across firm sizes, following market capitalization categories: micro-cap (¡\$300M), small-cap (\$300M–\$2B), mid-cap (\$2B–\$10B), large-cap (\$10B–\$200B), and mega-cap (¿\$200B). In addition, we incorporated firms from over 30 states and international jurisdictions to capture regional differences in reporting and tagging conventions. This diversity ensures that the benchmark reflects realistic heterogeneity observed in financial disclosures across sectors, scales, and regulatory environments.
Figure 5. Distribution of Tickers by Industry Sector.
Figure 6. Distribution of Tickers by Market Cap Size.
Figure 7. Distribution of Tickers by Geography.
Appendix E Benchmark Granularity Justification
Each 10-K report follows the SEC’s XBRL filing standard, where all narrative and tabular disclosures are linked through machine-readable tags defined in the US-GAAP taxonomy. While the XBRL framework provides document-level structure via instance, schema, and linkbase files, the tagging of financial values and concepts occurs at a much finer granularity—typically within individual sentences or table rows that contain numeric facts and their contextual descriptions.
For this reason, our benchmark adopts the sentence and table sequence as the fundamental input unit rather than entire reports. This design choice is motivated by three considerations:
  • Locality of Semantic Scope: Financial facts and their US-GAAP tags are usually expressed within a single sentence or a bounded table region. Modeling at this level aligns the input scope with the tagging scope, reducing noise from unrelated content within long filings.
  • Computational Efficiency: Full 10-K filings often exceed hundreds of pages and contain heterogeneous sections (e.g., risk factors, management discussion, and financial statements). Processing the entire report as one input sequence would exceed model context limits and obscure local numerical relations. Segment-level inputs enable fine-grained supervision and efficient evaluation.
  • Faithful Tag Alignment: Because each numerical entity is explicitly linked to its taxonomy tag in the XBRL instance document, using sentence or table segments ensures a one-to-one correspondence between input content and ground-truth annotations. This granularity preserves the original tagging fidelity of the filings.
Overall, this segmentation strategy allows our benchmark to remain consistent with how XBRL tagging operates in practice (on localized factual statements), while maintaining scalability and interpretability across large collections of filings.
Appendix F Validation Guideline
F.1. Task Definition
Given a context and an extracted triplet consisting of:
  • an entity,
  • its entity type,
  • a US-GAAP concept,
Your task is to determine whether the triplet is correct.
Make a binary decision:
  • 1 for correct,
  • 0 for incorrect.
F.2. Validation Procedure
Follow the rules below, the definitions of entity type are shown in Table 12:
  1. (1)
    Entity Check Determine whether the extracted entity is a numerical value.
    • If yes, proceed to Step 2.
    • If no, set is_correct = 0.
  2. (2)
    Entity Type Validation Verify whether the identified entity type is correct based on the context and the definitions in Table 12.
    • If yes, proceed to Step 3.
    • If no, set is_correct = 0.
  3. (3)
    US-GAAP Concept Validation Assess whether the assigned US-GAAP concept is appropriate based on the taxonomy tag definitions.
    • If yes, set is_correct = 1.
    • If no, set is_correct = 0.
Table 12. Entity type definitions used for validation.
Entity Type Definition
monetaryItemType Financial amounts expressed in currency, such as revenue, profit, or total assets.
integerItemType Counts of discrete items, such as the number of employees or total transactions.
perShareItemType Per-share values, such as earnings per share (EPS) or book value per share.
sharesItemType Counts of shares, such as outstanding shares or ownership stakes.
percentItemType Ratios or percentages, such as tax rates, growth rates, or discount rates, usually expressed with a percentage symbol (%).
F.3. Important Instructions
  1. (1)
    Non-Arabic Formats Financial numerical entities may appear in word form (e.g., ten million) and must be correctly identified and converted into standard numerical format.
  2. (2)
    Magnitude Terms If a number is followed by a magnitude term (e.g., hundred, million, billion), do not expand it into the full numerical value:
    • Two hundred $\rightarrow$ extract only two, not 200.
    • 10.6 million $\rightarrow$ extract only 10.6, not 10,600,000.
  3. (3)
    Standardization of Format Remove formatting symbols while preserving the numerical value:
    • Remove currency symbols (e.g., USD).
    • Remove percentage signs (e.g., %).
    • Remove commas (e.g., 1,000 $\rightarrow$ 1000).
Appendix G Evaluation Metrics
@@
To provide a fair evaluation of overall benchmark performance, we adopt a set of metrics, focusing primarily on macro-level and micro-level evaluation strategies inspired by the previous work (Sharma et al., 2023). Macro-level evaluation computes precision, recall, and F1 scores independently for each BIO-concept label derived from the US-GAAP taxonomy, and then averages them without weighting. This ensures that each concept, including rare or infrequent ones, contributes equally to the final score, making it especially suitable for domains with skewed label distributions. In contrast, micro-level evaluation aggregates token-level true positives, false positives, and false negatives across all labels before computing precision, recall, and F1. This approach emphasizes the model’s overall tagging accuracy by treating every token equally and thus better reflects performance on frequent concepts. Together, these two metrics provide a balanced view of both per-concept performance and overall tagging quality.
For the FinNI subtask, the objective is to extract correct (entity, type) pairs, that is $(\texttt{Fact},\texttt{Type})$, from the financial document. Let $\mathcal{G}={(e_{i},l_{i})}$ denote the set of ground-truth (entity, type) pairs, and $\mathcal{P}={(e^{\prime}_{i},l^{\prime}_{i})}$ denote the set of predicted (entity, type) pairs. We evaluate the performance based on the following metrics:
(5) $\text{Precision}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{P}|}$
(6) $\text{Recall}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{G}|}$
(7) $\text{F1}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
@@
(8) $\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\hat{c}_{i},c_{i}^{*})$
where $\delta(\hat{c}_{i},c_{i}^{*})=1$ if $\hat{c}_{i}=c_{i}^{*}$, and 0 otherwise.
Appendix H Evaluation Models Details